OpenStudy (anonymous):

Reciprocal sum is exactly 1 now

OpenStudy (anonymous):

Find the number of solutions to the equation\[\frac{ 1 }{ a }+\frac{ 1 }{ b }+\frac{ 1 }{ c }+\frac{ 1 }{ d }=1\]

OpenStudy (anonymous):

where a, b, c, d are positive integers and a≤b≤c≤d.

OpenStudy (anonymous):

so far i got bcd+acd+abd+abc=abcd

OpenStudy (anonymous):

since its 3 products 3^4=81